430 MHz Gregorian Dome Receiver Log

as of Friday, 03-Dec-2021 05:29:51 AST

Go to Most Recent Entry
ANS anolla@naic.edu 10/14/99 The dewar was warming up due to power failure.Start pumping it , it was recovered after 3 hours.Also the 7/8 coax line from chn. A was removed to be used at the CH 1 430 Mhz receivers ( CRR'S )
ANS anolla@naic.edu 10/15/99 The 7/8 " cable was hooked up back. But there is a problem with the tsys. ,it is twice as high as it was in that channel. Have to be recalibrated .
ANS anolla@naic.edu 11/29/99 Receiver turned off in order to replace A channel amplifier. It was found during the 11/25/99 maint. that the first stage of channel A was burned out.
ANS anolla@naic.edu 12/17/99 W Genter replaced the burned out (chnn. A) amp in the dewar. Now the s.n. for these amp is 108 and for chnn. B is 109.Also the protection ckt was removed in order to be check ,chnn. B isolation is 37 db and chnn. A is 42.2 db. It was found that protective device for chann. A had an intermittent contact,it was fixed. Bias for the amplifier S.N. 108: Stage 1 1.5V/5.0 mA, Stage 2: 2.4V/8.0 mA
ANS anolla@naic.edu 12/27/99 Protective devices were installed back , receiver was checked by Tel. Operations
ANS anolla@naic.edu 1/5/2000 The waveguide was flooded with water ,causing that chnn. B was with a Tsys. of 690 K deg. In other words , if you got chnn. A OK and chnn. B bad ,you got two conditions or a burned out amp. or sharks inside the waveguide (water)
ANS anolla@naic.edu 1/12/2000 Jose Vives installed a window in the horn in order to keep moisture out of the system.Also we recommend to pressurize it.
ANS anolla@naic.edu 1/17/2000 System warmed due to contamination, cryo crosshead turned off in order to replaced it.
ANS anolla@naic.edu 5/17/2000 Receiver Calibration Results. Channel A: Att (dB) / Cal Mode / % R/S / Cal Value (K) 29.53 Lcorcal 95.5 10.5 28.85 Lxcal 98.8 12.46 28.53 Luncorcal 96.9 10.68 28.85 L90cal 97.0 12.49 19.90 H90cal 103.9 105.10 19.90 Hxcal 100.6 101.74 19.90 Huncorcal 102.7 103.86 19.90 Hcorcal 100.3 101.43 Channel B: 29.53 Lcorcal 105.4 11.61 29.53 Lxcal 96.1 10.58 29.53 Luncorcal 96.7 10.65 29.53 L90cal 101.4 11.16 19.90 H90cal 89.5 90.48 20.67 H90cal 106.3 89.99 19.90 Hxcal 97.3 98.40 19.90 Huncorcal 89.5 90.48 19.90 Hcorcal 97.3 98.40
ANS anolla@naic.edu 6/1/2000 Because system was cycling cryo system was shutted off yesterday in order to warm it up and replace the crosshead.The crosshead & cryo lines were vac pumped and the vac drop to 4 un's. Then today we started the cooldown.
ANS anolla@naic.edu 6/6/2000 Cooldown was successfully , system is running normally .The 16 deg stage is around 12 to 14 deg. As always 70 deg stage is reading wrongly, instead around 70 deg is reading 22.5 deg.
ANS anolla@naic.edu 8/9/2000 Dewar was turned off due to WG installation.Also the crosshead has to be replaced due to contamination , was cycling.
ANS anolla@naic.edu 9/28/2000 Dewar off due to bad cryo compressor that was shorted out by an AC line inside of it.
ANS anolla@naic.edu 9/29/2000 The amp Berkshire sn 109 of chnn. B was burned out.Been sended to be repaired at Berkshire.The spare one is missing , is model U-440 sn 106 .Bershire amp model U-430-2 sn 102 is going to be used , it has 7 db less gain but is usable .(instead 37 db is 30 db)
ANS anolla@naic.edu 10/4/2000 Receiver was cooled down. Chnn. A T Sys is 70.9 K & chnn B is 64.7K
ANS anolla@naic.edu 10/9/2000 Receiver was warmed due to a power dip on Saturday that tripped off the compressor. Cooled down.
ANS anolla@naic.edu 10/25/2000 System warmed up to replace crosshead,it was highly contaminated.
ANS anolla@naic.edu 11/2/2000 Dewar removed from system in order to replace it with the new one.RF filters are in the receiver cabinet N , drawer C2.
ANS anolla@naic.edu 12/11/2000 New Dewar installed at the Dome on 11/21/2000 . The cryo is running with a 350 cryo comp. , also a new open frame cabinet was installed for the Bias & Control chassis. Bias are at 430 Mhz : Amp A1 Vd 1.51 V Id 3.3 ma , Amp A2 Vd 1.5 V Id 6.1 ma Amp B1 Vd 1.52 V Id 4.3 ma , Amp B2 Vd 1.5 V Id 6.2 ma
ANS anolla@naic.edu 12/12/2000 VI remove the Monoplexer from chnn. A because a crack on it due to the preassure of the 7/8" coax cable that goes to the Dewar, in order to fix it. A bracket was build in order to relief the strain at the monoplexer. Also during the monoplexer check out at the lab it blew up a fuse at its driver, after replacing it took fire and that's made feel chicken about these new mono drivers. If they are going to be problems seekers is to be worry, because everytime they fail is a burned amp.
ANS anolla@naic.edu 12/22/2000 VI installed back the monoplexer , but is not hooked into the recvr until the mono driver is installed.JR suggested to wire the chn A A/L relay as chn B .So when 430 xmmtter is on the air the Jennings relay on chn A automatically goes to load .
ANS anolla@naic.edu 12/26/2000 In order to protect the chn A from 430 xmmiter , the A/L sw was wired as chn B. There is a jumper between second wire of PR 76 & first wire of PR 77.
ANS anolla@naic.edu 12/27/2000 Receiver wired for R.A. only ( A/L Jennings Sw's)
ANS anolla@naic.edu 1/3/2001 The receiver Hi Cal was calibrated and is as follow : Channel A Hi 90 = 92.5 k Hi X = 94.5 k Hi 0 = 95.0 k Hi Corr = 95.99 k Channel B Hi 90 = 90.38 k HI X = 102.6 k Hi 0 = 91.68 k Hi Corr = 102.5 k Note: The diode en at 430 Mhz is 15.325 db and the 20 db pad value was 19.94 db. Also a 3 db pad was added to the recvr noise cal out at channel A in order to have around a 100 cal, and on channel B a 5 db att was added for the same purpose.
ANS anolla@naic.edu 1/4/2001 Low Cals are: Chann A 0=10.1 k X= 12.3 k 90= 12.57 k Corr = 9.95 k Chann B 0=11.96 k Uncorr X= 10.34 k 90= 11.85 k Corr = 10.4 k Notes: A/L chann A is 5.36 db , chann B is 5.73 db , T Sys both channels 95 k
ANS anolla@naic.edu 1/18/2001 The cryo system comp replaced , it was comp 008 and it was replaced with comp 012 , lines were 6 & 7 , now are 12 & 13 . And is running with CB Dewar.
ANS anolla@naic.edu 2/2/2001 Comp 012 replaced with comp 009 , it is running in pair with CB Dewar.
ANS anolla@naic.edu 2/6/2001 Monoplexer removed from receiver A , also the A/L sw interlock with Xmmter.
ANS anolla@naic.edu 3/6/2001 Dewar was warming up because the crosshead driver was unplugged accidentally. It was cooled down again.
ANS anolla@naic.edu 3/9/2001 The Dewar was removed in order to have the xmmter mode using a Lunar amp.
ANS anolla@naic.edu 5/17/2001 Xmmiter mode installed and works successfully. Went up to 1-megawatt .The Dewar receiver was used.
ANS anolla@naic.edu 5/21/2001 Working with the R.A./Radar transfer sw .EC wants that the these will be performed automatically and the only way is adding a sw to the LRR 2 transition that when it is removed to install the xmitter into that port will sw the LRR 1 A/L sw to the no interlock mode.
ANS anolla@naic.edu 5/23/2001 The A/L sw at radar room bypassed, recvr always on antenna, except when 430 MHz xmmter at CH 1 up.
ANS anolla@naic.edu 7/16/2001 Both rf amps at the Dewar were burned out during the dual beam run on 7/12/2001.They were replaced on 7/13/2001 , the sn for the new amps at the Dewar are: Chnn.A sn 450-12: Stage 1: 1.5V / 3.7 mA, STage 2: 1.52V / 5.7 mA Chann B sn 450-13: Stage 1: 1.5 V / 3.9 mA, Stage 2: 1.5V / 5.4 mA It is suspected that these amplifiers were burned due to a faulting timing box that was producing strange glitches on it.( Also the rf amplifiers of the CH 1 were burned out) The burned out ones were send to Richard Bradley at Charlottesville to be repair. They are Sn 450-10 and Sn 450-11.
ANS anolla@naic.edu 7/19/2001 An L.O. gating sw was intalled .It is between chann A 300 mhz L.O. (SW 4 L.O. distribution chassis UG801D10) and the converter chassis ( UG801D09 ).
ANS anolla@naic.edu 7/20/2001 The ant/load Jennings sw for channel A is completely burned out.It was replaced with the one on chnn B.There was a barely A/L change ,less than 2 db and it is supposed to be +/- 5 db.
ANS anolla@naic.edu 7/23/2001 System seted back to RA use. Chann B A/L Jennings replaced with an spare (the spare for CH 1 A/L sw).
ANS anolla@naic.edu 7/24/2001 System back to radar mode.
ANS anolla@naic.edu 7/26/2001 There was RF arking at the Dome reported .We check inspecting at the Dome but there was not any arking at all. But we think that there is something wrong yet.
ANS anolla@naic.edu 7/27/2001 The RF amp at chnn A was burned out, it had just 2 db A/L change and Tsys was 300 k. Replaced it with the chann B RF amp and it has a Tsys of 79 k.
ANS anolla@naic.edu 7/31/01 repaired amps 450-10 and 450-11 returned from r. Bradley in Charlottesville.
ANS anolla@naic.edu 7/30/2001 System turned off in order to replace burned RF amp in chann, A. It will be replaced with amp sn 450-14 and its biases are: STage 1: 1.5V / 3.6 mA, Stage 2: 1.5 V / 6.3 mA Also the A/L system was returned to RA mode from radar.
ANS anolla@naic.edu 8/1/2001 System was cooled down after replacing the chann A burned out amplifier. It was tested at room temperature because we were not able to move the rotary floor (L. Baker was taking data) to take a cooled Tsys.Also the supply cryo line filter was replaced with a serviced one (005).
ANS anolla@naic.edu 8/9/2001 The ED 009 cryo compressor was taken down for service, so comp ED 011 was hooked to the 430 mhz and the CB.
ANS anolla@naic.edu August 31, 2001 The ED 011 cryo comp was removed from system to hook the proper one to it, now sys has comp ED 012 on it. Also we suspect that comp ED 011 is leaking.
ANS anolla@naic.edu 9/13/2001 The chann A ampl was burned out during the dual beam run , biases looks ok but amp is dead. It was replaced with chann B one.
ANS anolla@naic.edu 9/17/2001 Chann. A amp replaced, it was burned during the dual beam run. The burned out sn is 450-14. It was replaced with amp sn 450-10 (spare one)
ANS anolla@naic.edu 10/10/2001 New monoplexer mounting bracket installed, it is more well design. Built by H. Monroig.
ANS anolla@naic.edu 10/20/2001 System installed for the dual-beam run.
ANS anolla@naic.edu 10/24/2001 System back to normal installation after a useful dual-beam run.
ANS anolla@naic.edu 10/24/2001 430 mhz A/L sw installed thru the CB/SBW recvr shutter control relay to protect it from the SB xmitter.
ANS anolla@naic.edu 11/13/2001 System installed for the dual-beam run.
ANS anolla@naic.edu 11/15/2001 System back to normal , successful run ,no burned amps these time. We just notice that the CB & WSB shutters were not closing and it is due to faulty software.
ANS anolla@naic.edu 12/11/2001 Channel A tsys is 90 ,Ganesh measured on 12/7/2001. So , EC told us to use chann B instead A for the dual-beam run on these week.
ANS anolla@naic.edu 12/13/2001 R.F. amp & monoplexer were burned yesterday afternoon , so this morning we put the system back to normal. Tomorrow we will replace both amps.
ANS anolla@naic.edu 12/14/2001 Both rf amps at the Dewar replaced ,were burned out. Chann A amp was sn 10 now is sn 11. Chann B was sn 13 , now is 12. I set the bias for chnn A that were Vd 1.51 vdc and Id 4.1 ma , no bias data on amp B.
ANS anolla@naic.edu 12/20/2001 There is still a problem with the Tsys of channel A, it measure 100k instead +/- 70k.
ANS anolla@naic.edu 12/21/2001 JR found water at the 7/8" elbow at the input of the Dewar on chan A. This was the problem with chan A high temp, now both chan are +/- 70 k.
ANS anolla@naic.edu 1/21/2002 System installed for dual-beam run.
ANS anolla@naic.edu 1/22/2002 System back to normal ( R.A.)
ANS anolla@naic.edu 1/24/2002 The RF gating sw was taken down in order to install in the same chassis post amplifiers to be used when receiver is in the RA mode; there is not enough gain in these mode.
ANS anolla@naic.edu 1/31/2002 There are 2 amplifiers that are been fixed at these moment , these are sn 450-13 and 450-14 ,but were burned out during dual beam runs . The 450-14 was burned during a run on Sept. 2001 and the 450-13 in Dec. 2001.The amplifiers belongs to a new project to replace the CH 1 Dewar , Ed Castro has them for these purpose.Richard Bradley is the contac person at factory.
ANS anolla@naic.edu 2/14/2002 System set for dual beam operation.Checking the set up was found again that the CB & WSB shutters were open , there is something wrong with the software commands.
ANS anolla@naic.edu 2/18/2002 System back to R.A. mode.
ANS anolla@naic.edu 3/25/2002 System Set for dual beam run , will be until 3/27/2002.
ANS anolla@naic.edu 3/28/2002 System set back to R.A. Channel A has a T sys of 200 k . We double check the rf installation and found plenty of water in the chann A elbow at the Dewar.Suspect that the amp has been degraded with the run doesn't matter if Vd & Id can be adjusted.
ANS anolla@naic.edu 3/29/2002 Amp in channel A replaced with the spare one ( sn 450-10 ,by the way no more spares , all at NRAO been fixed) . Burned one delivered to Lisa.
ANS anolla@naic.edu 4/1/2002 Both channels with a Tsys of +/- 60 k.
ANS anolla@naic.edu 4/15/2002 System set for dual beam , Tsys for channel A thru monoplexer was 70 k.
ANS anolla@naic.edu 4/17/2002 System at room temperature , cryocomp quit ( ED012) . They decided to keep running at room temp.
ANS anolla@naic.edu 4/19/2002 Amp 450-10 on chan A & 450-12 on B. Amps 450-13 & 14 came back from NRAO , were burned.
ANS anolla@naic.edu 4/22/2002 Channel A got burned out again ,it had a tsys of 100 k ,chann B was 60 k .We don't know what's going on with this problem , EC suspect that the power density in the rotary floor is too high .This is the second time in a roll that it happen. So Dewar was turned off in order to replace burned amp tomorrow. Also the cryocompressor was replaced with ED 007 , ED 012 quit working.
ANS anolla@naic.edu 2/24/2002 Channel A amp replaced ,it was burned out during the last dual beam run.Burned out one is sn 450-10 , replacement is 450-13. Tsys are chann A 57.8 k and chann B is 54.0 k. Lisa is going to send it for repair ,also the 450-11 that was previously burned
ANS anolla@naic.edu 5/13/2002 The RF & LO gating sw's were TTL isolated with an HP optocoupler .It was done in order to have the proper TTL level at the gating sw drive .
ANS anolla@naic.edu 5/16/2002 System replaced for dual beam.Chann A amp is at room temp. JR measured the output of the monoplexer and found that was 30 dbm level coming from it. After a trial run the room temp amp got burned , so they cancelled the dual beam run. Have to find out what is going at the Dome with the reflected power.
ANS anolla@naic.edu 5/20/2002 Amps 450-10 & 450-11 are with Richard Bradley in order to be fixed , they were burned out.
ANS anolla@naic.edu 6/11/2002 System sw to dual beam operation ( radar ). It's using a Lunar amp (room temp) with a Tsys of 120/170 k .
ANS anolla@naic.edu 6/13/2002 The Lunar amp was burned yesterday. So we replaced it with an spare one.Also we hooked up a Melabs circulator in the output of it for more isolation and replaced it's rf cable between it and the post amp.
ANS anolla@naic.edu 8/1/2002 There are 2 cooled spare amps for the receiver they are sn 450-10 & 450-11.
LSW lwray@naic.edu Wednesday, August 7, 2002 15:20:03 Changes to the front end of the receiver: More gain for radio astronomy and radar modes and more protection against high power signals into the amps. Details: Added a limiter (limits power to LNA), new (shorter) cable from circulator to gating switch, completed new gating switch interface, included additional gain from single stage PHEMT LNA 430 GSE after the gating switch.
LSWlwray@naic.eduThursday, August 8, 2002 16:01:33received 3 new hi-level PHEMT (HY430GSE-34 amps from Angle Linear, S/N# 11417-11419, plus the repaired #11273.
ansanolla@naic.eduFriday, August 9, 2002 15:00:01System installed for dual-beam run. Using the Angle Linear amp (room temp) it has a Tsys of 170.
ansanolla@naic.eduMonday, August 26, 2002 10:18:20System was returned to RA on 8/15/2002.Forgot to write it up in log book.
LSWlwray@naic.eduWednesday, September 4, 2002 16:30:55Post amp chassis (first post amp, the one with the gating switch) was taken down for the day, to switch the low-level Angle Linear amp with a high-level amp. The Radar front end, box #1 had the high-level amp switched with the low-level.
ansanolla@naic.eduWednesday, September 25, 2002 11:50:03Cryocomp shut off due to power dip. Dewar warmed up , been cooled. NO REMOTE SW ON IT.
aansanolla@naic.eduWednesday, October 9, 2002 12:41:41The Angle Linear low level amp was burned out , it was replaced with the spare high level one (amp 3).There is an spare amp at the Dome and it is at the Gray box at the floor.
ans anolla@naic.eduWednesday, October 9, 2002 12:44:25The Dewar vacuum was bad making cryotemps on it beeing high.Pumping it solve the problem.
ansanolla@naic edu.Thursday, October 10, 2002 13:58:51The monoplexer got burned out. LSW checked all the amps ( 3 ,CH 1, Dome & spare).They were ok.Also mono at CH 1 got burned.
LSWlwray@naic.eduWednesday, October 16, 2002 11:39:49Found cable was bad (centre pin too short) on cable from Ch A post amp to IF/LO cabinet. The power levels at the upstairs power meter showed 20 dB less power than Ch B. This explains the variable gains seen during the world day radar run. The cable is being repaired.
LSWlwray@naic.eduThursday, October 24, 2002 08:22:13Switched bandpass filter 425-435 to 408-412 on October 23. Plan to switch back to normal on October 27.
ansanolla@naic edu.Monday, November 11, 2002 12:41:35Dual beam run on the way , Angle Linear amps on (Dome & CH 1). Dome Tsys was 148k .
ansanolla@naic.eduFriday, November 15, 2002 13:12:10System back to RA due that a ligthning struck the power lines killing the power supply to the site.
ansanolla@naic.eduTuesday, November 26, 2002 11:27:42System set for dual beam run. Tsys was 145k. Angle Linear amp # 2 with it.
ans anolla@naic edu.Wednesday, November 27, 2002 13:14:59System back to RA - cooled receiver, monoplexer removed. Tsys for both channels is 65k.
ansanolla@naic edu.Monday, December 2, 2002 13:54:47Syst for dual beam run. Tsys was 150k at room temp.
ansanolla@naic edu.Thursday, December 5, 2002 15:28:39Sytem back to RA , dual beam amp downstairs to be checked.
ans anolla@naic edu.Monday, February 3, 2003 13:23:23System back to RA , it was at the Radar mode (Angle linear amps )since 1/28/2003.
ansanolla@naic.eduTuesday, May 20, 2003 13:10:14The post amps filters (430 mhz , 10 mhz BW ) were removed for a T. Goshi observation.
lswlwray@naic.eduSunday, May 25, 2003 15:21:29May 21: 10 MHz filters put back in, system back as normal.
ansanolla@naic.eduTuesday, May 27, 2003 15:19:33System was set for dual beam run.It was using the cooled amp ,but due to a crowbar the cooled amp was burned , replace it with the Angle Linear room temp amp.
ansanolla@naic.eduMonday, June 16, 2003 14:31:53Due to intermod in the system as in the CH #1 , the room temp amps were installed on the system to avoid these intermod , but somebody reject these because they need the cooled amps in the system.
ansanolla@naic.eduWednesday, June 18, 2003 09:44:09During the last dual beam run on 5/27/2003 tha amp 450-13 was burned out due to a crowbar.Delivered to LSW in order to be repaired at factory
lswlwray@naic.eduMonday, June 23, 2003 09:02:53May 29: after last radar run, #450-13 in Ch A was replaced with #450-14.
ansanolla@naic.eduTuesday, June 24, 2003 13:05:15System was set for dual beam run.It's Tsystem is 150K ,Ganesh suspect that the monoplexer driver is not working properly giving us such Tsys , he was expecting a Tsys of 100/130k.
ansanolla@naic.eduThursday, June 26, 2003 15:15:29System back to R.A.
ansanolla@naic.eduTuesday, July 1, 2003 10:16:15System set for dual beam run until next Monday. Ganesh tuned the amp and the monoplexer in order to obtain a better Tsys , rigth now is less than 100k
lswlwray@naic.eduTuesday, July 1, 2003 12:42:42Amplifier 450-13 returned from NRAO, repaired. Bias settings tested: 2V/10mA, room temp performance ~18 dB gain, ~54K noise temp.
ansanolla@naic.eduWednesday, July 9, 2003 12:48:47Channel A had a higher Tsys than chann B. We found that the inner 7/8" coax connector at the Dewar was loose ,was tightened and Tsys returned to normal.
ansanolla@naic.eduWednesday, July 23, 2003 09:10:27System set for dual beam run. Tsys with Anglelinear amp was 101 k.
ansanolla@naic.eduThursday, July 24, 2003 15:25:39There was a complaint that some rf was getting into the Dome amp.Found that the Dome monoplexer had leaky diodes, so the CH 1 monoplexer was swap to the Dome ( 8 diodes on it) and the spare one was installed in CH 1 ( 6 diodes one).
ansanolla@naic.eduMonday, July 28, 2003 15:49:38System set for RA. Its Tsys are chann A 61.7k and B 62.3k.
ansanolla@naic.eduFriday, August 1, 2003 14:06:50The 430 mhz rf filters 10 mhz BW were replaced by a new ones but with a BW of 20 mhz , also Ganesh choose a couple of filters to act as rejection ones to avoid TV chann 22 rfi.
ansanolla@naic.eduMonday, September 22, 2003 11:16:20System set for dual beam. Tsys was a 100k.
ansanolla@naic.eduMonday, September 29, 2003 16:04:17System back to R.A.
ansanolla@naic.eduWednesday, October 8, 2003 12:47:04Channel A had a Tsys 30 higher than channel B. As always the contac of the RF cables after the dual beam run is hurting the system badly.We clean them up and normal T sys came back and that's bad.
ansanolla@naic.eduMonday, October 20, 2003 14:01:59The new shutter driver at the Dome was producing acomb generated rfi to all low frequency receivers at the Dome(327,430,611,etc.)
ansanolla@naic.eduMonday, October 20, 2003 14:36:54While replacing a burned out Angle Linear amp (low level , model 430GST )from the last dual beam run , I found that the sn from the replacement one and the burned one had the same sn , 12961 ,so I add an -1 ID to the bad one and a -2 to the good one.
ansanolla@naic.eduTuesday, October 21, 2003 13:18:48System is in the dual beam run. Tsys is a 100k . Spare amp on top of waveguide plus plier .
ansanolla@naic.eduFriday, October 24, 2003 14:21:47Dome 430 mhz system back to radioastronomy.Tsys at chann A for radar was with Angle Linear amp 81k at room temp.The RA Tsys is for chan A is 50.4 and for chan B is 48.6.
lswlwray@naic.eduThursday, October 30, 2003 16:07:36Antenna-Load high power Jennings relay had extensive carbon deposits on the Antenna port of Ch A. Caused high Tsys in Ch A . Replaced with spare, re-wiring of connector needed to be done (the connector isn't exactly the same). Also carbon found in connector to waveguide pickup. This was cleaned and replaced. Tsys now ~65K for Ch A and ~60 for Ch B.
ansanolla@naic.eduTuesday, November 11, 2003 12:54:26System set up for dualbeam run. Tsys was 108k with amp # 3 ,low level one.Spare one is amp # 6 , a hi level one.
ansanolla@naic.eduWednesday, November 12, 2003 09:53:24The A/L relay of channel A( Jennings Sw ) got bad , it started blowing the fuse at the Jennings P.S. It was replaced by the one in channel B. There is not an spare one.
ansanolla@naic.eduMonday, November 17, 2003 15:48:36System back to R.A. , channel B A/L sw is a Dow-Key relay. The Jennings relay at channel A got bad , so we took the channel B one and installed it on A.The only replacement that we found was these Dow-Key supplied by Iguina.
ansanolla@naic.eduTuesday, December 16, 2003 14:18:08System set for dual beam . Angle Linear amp #3 on it ,Tsys is 123k.There are two spare amps at the Dome ,amp #4 (low level) and amp #6 (high level).
ansanolla@naic.eduWednesday, December 17, 2003 13:32:31System back to R.A. due to 430 mhz xmmiter failure. One klystron quit working.
ansanolla@naic.eduThursday, December 18, 2003 13:01:03The RF gating sw for dual beam run was replaced with a Mini-Ckts model ZASWA-2-50DR , it has an isolation of a 100 db instead 50 as model ZYSWA-2-50DR.
ansanolla@naic.eduWednesday, January 7, 2004 15:06:33The cryocomp ED007 was replaced by ED011 due to an oil leak on it.
ansanolla@naic.eduWednesday, January 14, 2004 13:01:22On 1/8/2004 the cryocomp was replced again due that the ED011 quit working after less than 24 hours.But on 1/9/2004 we found that the ALFA cryo comp used to replace the ED011 was contaminated , making that both cryopumps (430 & CB ) got stock.After purging the CB one we started it's cool down.430 mhz cryo pump was atached to cryocomp ED008 after been purged.
ansanolla@naic.eduTuesday, January 20, 2004 16:01:20System cryo , shut off in order to replace cryopump that got contaminated by the ALFA receiver. Was ED018 , is ED025.
ansanolla@naic.eduThursday, January 22, 2004 12:39:30System was found after cool down with it's Dewar vacuum lost , it was due to it's pup up valve . Was sealed with vacuum seal and it cooled again.
lswlwray@naic.eduTuesday, March 30, 2004 10:11:20amplifier 450-12 sent to NRAO for repair.
ansanolla@naic.eduMonday, April 5, 2004 11:00:25Dual beam run since 3/29/2004 , Tsys cooled chn A 64k , chan B 75 k. With room temp was 45k ??????. System returned to normal today. It was not possible to take it's Tsys.
ansanolla@naic.eduMonday, May 17, 2004 12:48:29Dual beam run since 5/17/2004 up to 5/20/2004. Dome Tsys was 100k with room temp amp # 3 (low level). The spare one is # 6 and is a high level one (there is another one at the plastic box in the corner , is #4 and is a low level)
boobooTuesday, May 25, 2004 10:21:57boo
lswlwray@naic.eduTuesday, June 8, 2004 11:53:20Current amplifier S/Ns in cryo dewar are: Ch A: 450-14, Ch B: 450-12
lswlwray@naic.eduTuesday, June 15, 2004 08:18:48June 14: Dual beam installation day. Using room temperature amplifier in dome. Cryogenics amps are turned off, so TsysA: room temp amp, TsysB: turned off cryo amp.
lswlwray@naic.eduMonday, June 21, 2004 08:40:04June 18: Dual beam removal today.
ansanolla@naic.eduTuesday, August 17, 2004 09:19:20System set to dual beam operation on 8/16/2004 AST 11:30.
ansanolla@naic.eduTuesday, August 17, 2004 12:42:12Regarding last entry: Tsys was measured to be 105K.
ansanolla@naic.eduThursday, August 19, 2004 13:32:48System set to R.A. from dual beam. Tsys is chan A=68 deg K chan B=78 deg K .
ansanolla@naic.eduWednesday, August 25, 2004 11:41:48The 430 & CB receivers cryocomp was replaced , it was ED 007 , now is ED 012. The ED down for prev. maint. & modification.
ansanolla@naic.eduThursday, September 9, 2004 09:16:35Due to power breakdown of 1.5 hrs. system was warmed , but due to priorities it was left warm until today .
ansanolla@naic.eduFriday, September 10, 2004 11:03:40System set for dual beam run from 9/10/2004 to 9/16/2004. Tsys was for cooled amps ; A=55 deg & B=64 deg. For the dual beam Tsdys (room temp amp) is 112 deg (LL Angle Linear # 4).
ansanolla@naic.eduFriday, September 24, 2004 11:50:52Cavity filter with notch ones installed in channel A.Under test.Tsys was with it 55 k deg.
ansanolla@naic.eduMonday, October 11, 2004 12:39:59System was down due to power failure on Saturday , was cooled down by BG and Ganesh.
ansanolla@naic.eduTuesday, November 9, 2004 10:54:40System in Dual Beam set up.
ansanolla@naic.eduTuesday, November 9, 2004 11:48:18Forgot to mention that Tsys at room temp is 117 k.
ansanolla@naic.eduMonday, November 15, 2004 12:59:58System back to normal , it was dual beam. Tsys are chan A 73.1 k and B 74.3 k.
ansanolla@naic.eduMonday, December 6, 2004 14:45:23System in the dual beam mode.Tsys for Dewar was chann A 62& B 63. Tsys for the room temp been used with the dual been is 104k.
ansanolla@naic.eduFriday, December 10, 2004 15:50:29System back to normal. It was set for dual beam (tsys was 143k).Tsys for RA is chann A=68k & B=72k.
ansanolla@naic.eduFriday, December 10, 2004 15:52:59Forgot to mention that waveguide & transition is full of water due to condentation.The window for these receiver is broken , need replacemnt.
ansanolla@naic.eduThursday, January 13, 2005 14:05:42AC line cord guard added to the post amp chassis.
ansanolla@naic.eduTuesday, January 25, 2005 12:48:38The noise injection board at the post amp chassis replaced with the modified one. Is in order to avoid oscillation sources like the 60 Mhz in the SBL recvr.
lswlwray@naic.eduThursday, January 27, 2005 12:39:14Jan 26: installed DCI filter (bandpass + 4 notches at TV stations) on Ch A preceding the LNA. This is to reduce the intermods that occur inside the 430 band from the TV stations. Also Carlos moved the isolators inside the Gating switch box. The gating switch box now contains: switch, isolator, 375/250 MHz filter, Angle Linear postamp. After the gating switch the signal goes to the MIT filter and then postamp box.
lswlwray@naic.eduThursday, January 27, 2005 12:40:23Today: Nolla and Cheo installed a DCI (bandpass + 4 notches) filter on Ch B, after the cal coupler and before the LNA.
ansanolla@naic.eduWednesday, February 2, 2005 14:08:52On the week of Jan 18/21 the A/L sw was swap between CH 1 channel 2 and Dome channel 2. At the Dome the Dow-Key relay was replaced with the Jennings relay from the CH 1 chan.2 and vise versa at the CH 1 chann. 2. There was an improvement of Tsys at the Dome of 10 k better , it was a Tsys similar to channel 1.At the CH 1 the A/L was 7db but for some reason it was improved to 9db and Tsys was 40/45 k instead 60/65 k. Why ????
ansanolla@naic.eduTuesday, February 22, 2005 16:22:21There it was a power shut down at the Dome due to a braker shut down at the substation that supplies the power to it.It affected several receivers because it last more than 30 minutes.The recvrs were 430 ,SBH,SBL and SBN.Vacuum pump on all of them.These happened on 2/18/2005.
ansanolla@naic.eduFriday, February 25, 2005 16:23:00Jennings a/l relay from channel b removed in order to installed on the CRR/LRR 1 system at the C.H. 1.(the one there is not working properly).
ansanolla@naic.eduMonday, March 7, 2005 13:01:55Monoplexer installed at channel A for a test. Channel B at load all the time due to the removal of the interlock sw at wave guide.
ansanolla@naic.eduMonday, March 14, 2005 09:42:54System been running as dual.Room temp amp , Tsys too high.!!!! a 160k.
ansanolla@naic.eduTuesday, March 15, 2005 12:44:27System back to R.A. The dual beam was cancelled due to problems with the 430 mhz xmitter.
ansanolla@naic.eduMonday, March 21, 2005 15:21:59Cryocomp replaced , it had a knocking sound.It was ED 012 , now is ED 007.
ansanolla@naic.eduMonday, March 28, 2005 11:54:43Dual beam installed today.New Jennings sw at chann.A
ansanolla@naic.eduMonday, March 28, 2005 13:58:47The 430 Mhz xmitter blew up the klystron that was recetly replaced.
ansanolla@naic.eduMonday, March 28, 2005 14:03:22The CH #1 log in not working.
ansanolla@naic.eduThursday, March 31, 2005 12:49:40Dual beam cancelled due to 430 Mhz xmitter problems (bad klystrons).Back to RA . Tsys A= 55.7 & B= 52.3.
ansanolla@naic.eduThursday, March 31, 2005 12:51:53Forgot to mention that the A/L relay for channel B is a Jennings sw , so Tsys is +- 10 lower.
BJwgenter@naic.eduWednesday, May 11, 2005 16:55:09Receiver warmed up today due to compressor failure in the ALFA receiver.
ansanolla@naic.eduThursday, May 12, 2005 15:18:37System back to it's own cryo comp ED 007.Cryo temps were 10.4 & 36.9 , it's VPG 0 & vac 0 un;s.Comp press were 70/280 psi. T sys were A=68 k & B=69 k.
W. Genterwgenter@naic.eduMonday, June 27, 2005 11:12:01Hi level room temperature amplifier has been installed in CRR1 with limiter. Dewar bias turned off and Dewar input terminated.
W. Genterwgenter@naic.eduMonday, June 27, 2005 11:13:46Todays note continued: Dewar temperature readings will not be available until the bias supply is turned on.
ansanolla@naic.eduWednesday, June 29, 2005 08:53:59The last two entries by WG were for the CH #1 430 Mhz receiver.JRV was requested by Phil Perillat to use the IF gating sw in series with the HP frequency synth.
W. Genterwgenter@naic.eduMonday, July 18, 2005 14:14:00Horn cover broken. With the absence of a cover condensation can accumulate in the waveguide. Victor I. is looking into purchasing new cover material.
ansanolla@naic.eduWednesday, August 10, 2005 16:02:24System on dual beam operation. Tsys was 100k with room temp amp #4 & alaskan limiter . Spare amp next to monoplexer in plastic box.
ansanolla@naic.eduTuesday, August 16, 2005 08:50:54System back to normal.There it was a new set up for the LO gating and is was that since now on the HP synth out will be gated thru the LO gating sw , so all converters will be gated.
ansanolla@naic.eduFriday, August 26, 2005 13:01:04Due to power failure on 8/25/2005 system warmed up.Cooled.
ansanolla@naic.eduWednesday, September 7, 2005 14:53:21System set for dual beam run.Using cooled amps with the fast cooled Dewar.Tsys was 70K.
ansanolla@naic.eduTuesday, September 13, 2005 08:13:28System back to RA.Was with dual beam using the mini Dewar (fast Dewar).Tsys was with mini Dewar70k.
ansanolla@naic.eduWednesday, September 21, 2005 11:33:22System set for dual beam using the mini Dewar.It takes an extra hour of installation when running with the mini Dewar.
ansanolla@naic.eduMonday, September 26, 2005 08:57:20On 9/24/2005 the mini Dewar power supply blown its fuse letting the receiver to warm up. Bill Genter & Ganesh came to solve the problem.Also they noticed that there is not a power indicator in the P.S. unit nor label identifiying it.Going to take pictures of its location as reference for the future.
ansanolla@naic.eduMonday, October 3, 2005 09:43:20System in the dual beam mode. Tsys at Dome is 79.5k.Mini Dewar on it.
ansanolla@naic.eduMonday, October 10, 2005 15:06:04System back to normal.
ansanolla@naic.eduMonday, October 17, 2005 11:41:14The system waveguide coupler removed to measure it , Joe Greene in charge of it.System down until further notice.
ansanolla@naic.eduTuesday, October 18, 2005 15:33:02Dome 430 Mhz receiver back , but the D.C. wave guide still out until tomorrow. Tsys was chan A 78k & B 79k.
ansanolla@naic.eduMonday, November 7, 2005 14:11:24Channel B from mini Dewar is kaput , burned out. Also the Alaskan Limiter looks to be burned out ,too!!!!!!!
ansanolla@naic.eduFriday, November 11, 2005 11:08:44System back to R.A.
ansanolla@naic.eduWednesday, November 16, 2005 11:17:07Horn rexolite cover falling off.Waiting to recovered it at the Bowl?
ansanolla@naic.eduThursday, November 17, 2005 11:49:02Dual beam back. There was a problem of unstability in the system.it was caused by the monoplexer.It was the 16 msec pulse for gating the rf.Was causing reflections to the amp when gating.
ansanolla@naic.eduWednesday, November 30, 2005 13:25:00New braket installed to the 430 mhz Dome Jennings sw in order to reliev its weigth.
ansanolla@naic.eduThursday, December 1, 2005 08:55:52Also forgot to mention on yesterday log in that the xmitter crew replaced the rexolite Horn cover.
ansanolla@naic.eduTuesday, December 27, 2005 11:07:39Dual beam run until 1/3/2006.
ansanolla@naic.eduTuesday, January 3, 2006 11:55:33Dual beam removed.Mini Dewar was hot , its 10 amp cooling PS fuse was blown.Was cooled for test purpose and after half hour its Tsys was 87 K deg.The normal Dewar Tsys was 60 k deg for both channels.
ansanolla@naic.eduTuesday, April 4, 2006 09:19:32System was in dual beam operation since 3/29/2006.It was removed yesterday in the afternoon by Ganesh, Bill & Carlitos.It was removed 20 hours ahead of schedule due that power was run with the turbines.
ansanolla@naic.eduTuesday, April 4, 2006 11:36:22The local remote cal mode sw at the post amp replaced , it was a toggle one , was replaced with a push bottom one.In order to avoid mistakes.
ansanolla@naic.eduTuesday, May 9, 2006 14:33:27Cryo comp replaced , it was ED 007 now is ED 015.
ansanolla@naic.eduTuesday, May 23, 2006 13:25:02Cal chassis taken down to pinpoint a problem with the calibation of channel A diode,
ansanolla@naic.eduThursday, May 25, 2006 15:59:20JRV replaced the Cal TTl driver in the post amp chassis.He removed the inverter board of channel B.Also when installing the unit back found thar the Dome receivers cal driver unit had a bad power supply.
ansanolla@naic.eduWednesday, May 31, 2006 15:31:50Dual beam been run from 5/27/2006 until 5/31/2006.
ansanolla@naic.eduThursday, June 15, 2006 13:35:07System shut down for maintenance since 6/14/2006 AST 04:00.Cross head replaced at the Dewar it was ED 025 , now is ED 007.Also we found a vacuum leak at the Dewar , for some reason the pup-up valve on it is some mils smaller than it should be and when you remove the Dewar for cross head service any movement of it affect the vacuum creating a leak.
ansanolla@naic.eduFriday, June 16, 2006 10:58:14System up again since yesterday afternoon.
ansanolla@naic.eduWednesday, June 21, 2006 09:55:59calibrator taken down to track a big spike produced when firing cal.
ansanolla@naic.eduMonday, June 26, 2006 11:19:07Dual beam up until 7/5/2006.
ansanolla@naic.eduWednesday, June 28, 2006 09:32:03Dual beam run interrupted due to the replacement of a leaking waveguide section in the shielded room yard.
ansanolla@naic.eduFriday, June 30, 2006 12:55:42The Mini Dewar post amp (amp A) got burned out.It was replaced with our normal (R.A.)post amp.
ansanolla@naic.eduWednesday, July 5, 2006 15:44:12Dome 430 mhz receiver back to R.A. / Mini Dewar been taken down to repair the channel A that is burned.
ansanolla@naic.eduWednesday, July 12, 2006 11:52:01The Mini Dewar taken down for repairs , channel A burned out during last dual beam run.
ans anolla@naic.eduTuesday, July 18, 2006 11:49:04New turnstile installed , Tsys is chann A 55 K deg ,B is also 55 K deg.
ansanolla@naic.eduMonday, July 31, 2006 16:19:36No way to replace the amp chip for the mini Dewar , the burned out one in channel A. Rios was not able to desolder the burn out one in order to replace with the new one.Bill to order new ones.
ASRasantoni@naic.eduFriday, August 18, 2006 15:27:05During this week the removed jenning switch # 7247 ser # 749-1 was refurbish and tested by Alfredo Santoni. This one is keep it as a spare by Antonio Nolla of the Receiver Section.
ansanolla@naic.eduThursday, August 24, 2006 16:19:10The 430 Mhz Dome antenna cover was removed due that it make the isolation in the transition get worse.It was -60 db and it make it to -30db.So no cover for a while.
ansanolla@naic.eduFriday, August 25, 2006 10:18:09430 Mhz Dual beam on the run.This time the normal dewar being used ,using a new limiter on it. So we got these set up on system ; A/L sw , Cal D.C. , monoplexer,notch filter,limiter and then cooled Dewar.The mini Dewar is upstairs in case of an emergency.
ansanolla@naic.eduFriday, August 25, 2006 10:24:36By the way I forgot to mention that the mini dewar burn amp A was replaced by a new one.We send for two new ones but one got burned during installation, so no one left.
ansanolla@naic.eduThursday, September 7, 2006 11:35:47The ACLM limiter used during dual beam runs was brougth down to check it with the proposed mode sw (RA/Radar mode).
ansanolla@naic.eduWednesday, September 20, 2006 13:10:34Dual beam on until next week.
ansanolla@naic.eduTuesday, September 26, 2006 10:50:13Dual beam off.Systems ok.
ansanolla@naic.eduThursday, September 28, 2006 13:04:48Filter bank position 1 has new filters on it.They were fo 408 Mhz Bw 4 Mhz , now they are fo 430 Mhz BW 20 Mhz.
ansanolla@naic.eduWednesday, October 18, 2006 16:19:42There is a new RF/LO gating drive chassis in the 430 mhz system for radar run.It will be driven from gating signal coming from downstairs thru CH 1 gating chassis (f.o. connection from CH 1).It has 4 outputs one for RF ,one for LO & two spares.
ansanolla@naic.eduFriday, October 20, 2006 11:42:41Dual beam on , new RF/LO gating sw system.Gating pulse from downstairs thru CH 1.
ansanolla@naic.eduWednesday, November 15, 2006 16:00:51System down until tomorrow.Installing the new R.A./ Radar sw on it.
ansanolla@naic.eduThursday, November 16, 2006 13:00:28System back to normal.Now you can choose between RA or Radar mode for Dual beam run using a coax sw for these purpose.
ansanolla@naic.eduFriday, November 17, 2006 11:26:24System in dual beam mode.First time using the RA/Radar Mode Sw.
ansanolla@naic.eduTuesday, December 5, 2006 15:39:40Dual beam operation was removed since 11/22/2006.On RA mode.
ansanolla@naic.eduMonday, December 11, 2006 13:00:06JVR relocated the RA/Radar sw ,so in order to reach the amphenol connector to wire the Xmitter interlock sw.
ansanolla@naic.eduFriday, December 15, 2006 10:41:27System on dual beam operation. Tsys for chnn A 82.3 K deg.
ansanolla@naic.eduMonday, January 15, 2007 16:18:49Dome 430 Mhz channel A Jennings A/L sw replaced.Channel A was a little big higher on its Tsys than channel B (+/- 10 to 15 K deg) , after troubleshooting it pinpoint to the A/L sw.
ansanolla@naic.eduTuesday, January 16, 2007 11:25:10Cryo temps were high , 20 & 80 , vpg was reading 15. CR hooked up the vac pump and pumped the dewar for one hour , temps & vpg went down to 11.8 , 54.2 and vpg to 1.An eye on it.Valve closed.
ansanolla@naic.eduMonday, January 22, 2007 11:40:38System was set for dual beam run on 1/20/2007 at AST 10:00. It was supposed to run until 1/23/2007 AST 08:00 , but the 430 mhz xmitter developed problems and the system was returned to RA from radar during the morning time of 1/22/2007.
ansanolla@naic.eduTuesday, January 23, 2007 11:53:07BG took some data with the R & S Spectrum Analyzer to the receiver in the broad band mode . Looking for some signal that is saturating channel A at the Dome.
ansanolla@naic.eduMonday, January 29, 2007 11:24:47Dome 430 Mhz channel A was getting flucuations of more than 10 k deg higher on Tsys than channel B.Main suspect was the connection of the Jennings A/L/ sw and the Radar / RA mode sw (including the cal D.C.). So we cleaned up all the connections and tighth the plungers.System after that returned to normal . Tsys readings ,less than 2 K deg difference now.
ansanolla@naic.eduTuesday, February 13, 2007 08:01:28Forgot to mention that around 3 weeks ago we pumped the Dewar , its cryo temp started increasing.So it returned to normal cryo temp readings , but on thursday last week its cryo temp started increasing again ,so we pumped its Dewar again due that receiver was being used constanly.But during the week end started going up again ,so Ganesh told us to warmed it up and look for the leak on Dewar,walla we found the pup-up valve lay down in the floor and the Dewar with ice on it.Dewar was brougth down for check up.
ansanolla@naic.eduTuesday, February 20, 2007 15:02:36System back to normal on 2/16/2007. Bad pop-up valve seal at Dewar , the groove for the "O" ring was to deep.Replaced with tha 327 MHz Dewar one.
ansanolla@naic.eduMonday, March 19, 2007 11:50:27Dual beam run for a day.
ansanolla@naic.eduMonday, March 19, 2007 15:41:19System back to RA.It was found that the settings for the monoplexer drivers at the CH 1 & Dome were in the wrong position. Their setting have to be as follow for dual beam run , CH as master and Dome as slave.
ansanolla@naic.eduFriday, March 23, 2007 08:43:42Dual beam enable , until 3/26/2007 AST 16:00 hours.
ansanolla@naic.eduMonday, March 26, 2007 16:58:27System back to Radio Astronomy mode.
ansanolla@naic.eduMonday, April 23, 2007 09:42:10System down until further notice due to painting project.
ansanolla@naic.eduMonday, April 23, 2007 09:43:45System down until further notice due to painting project.
ansanolla@naic.eduMonday, April 23, 2007 16:22:48For special request from Alex we cool down system again.
CRVThursday, January 3, 2008 09:38:12 Vacuum pump connected to dewar on Dic. 7th,2007
CRVThursday, January 3, 2008 09:39:30Dewar cooldown started on Dic. 9th,2007 and it was cold in the afternoon.
crvTuesday, January 29, 2008 16:49:46The system warmed down during this weekend 430mhz dual beam run. The compressor braeker opened during thge power dips.
crvTuesday, January 29, 2008 16:51:46The system was vac. and cooled yesterday 1/28/08
crvMonday, May 5, 2008 10:36:05We placed a limitter at CH A postamp output.
crvWednesday, May 7, 2008 09:25:15Limitter change to the postamp. CHA output was actualized in the auto cad drawing.
crvMonday, April 13, 2009 16:49:12Filters on rcv. where removed to be send to calibrate and temporary replaced by 20MHz cavity filters
DEWdwhitlow@naic.eduWednesday, April 22, 2009 07:52:32Replaced CHA cooled LNA with uncooled LNA in hopes of improving recovery effect following each 430 radar TX pulse. Change was made on 21 April 2009 approx 1500 AST. Also measured monoplexer diode leakage current in RX state- result was 0.58 uA (at ~138 V bias). Also measured monoplexer spikes level at various places. While much larger straight out of the monoplexer, the peak level found was sub-mV at the output of the uncooled LNA, whose gain is ~20 dB. Evidently the spikes seen directly at the monoplexer output contain very little energy in the vicinity of 430 MHz.
crv Friday, August 14, 2009 16:13:00Dewar cryo turned off dewar warming up with vac. pump on to remove moisture from dewar.
crvThursday, August 20, 2009 14:57:44system started cooldown on tuesday august 18 2009
crvThursday, August 20, 2009 14:58:21system cold and vac. pump removed on wednesday august 19 2009.
crvThursday, March 25, 2010 16:28:51System started warming up on march 24th,2010.due to cryo pump problems. System is totaly warm doing vacc. today and the cryo pump will be installed march 26th,2010.
crvcrios@naic.eduFriday, March 26, 2010 13:40:03The cryo pump was replaced and is cooling. Cooldown started at 10:30Am and is reading 12.6 at the 16K stage at 1:40 Pm.
criosFriday, April 30, 2010 17:19:20System was warmed during the week because of racheting sound. Reciever started cooling yesterday on the new comp. C3 that was raise to replace the alfa compressor C2 tha failed.
criosWednesday, June 2, 2010 10:34:33System was moved to comp ed013 after replacing the compressor absorver on May 28th,2010.
crvWednesday, February 2, 2011 15:08:33there was problem with the ant/load. chassis R166 was removed from the control room and was cleaned system working.
crioscrios@naic.eduTuesday, May 10, 2011 16:06:29Channel A andrew connectors were retighten and Ch B Post amp output Cable was repaired.
crios Monday, June 6, 2011 14:03:10tsys on ch A were bad neg values. the problem was on chasis R166 on the control room. IC 9,10,11,13 were replaced (mc3487) tsys are ok.
Dana W. & Denis Urbaindwhitlow@naic.eduThursday, July 28, 2011 17:08:30Performed another round of Andrews connector tightening between WG-Coax transition and the dewar input. Monitored this activity with VNA at high display sensitivity (.05 dB/div) and noted that effects of pushing an shoving on fittings did make easily-discernible changes to transmission over frequency. We also found that one of the threaded studs where one of the Andrews lines goes into the DCI filter body has stripped threads in the soft aluminum filter body. Denis added a C-clamp to substitute for missing stud force. Did we do any good? Don't know yet- time will probably tell. We also have developed a fairly robust probe which allows injecting a signal directly into the horn. We ended up not relying on it today, but it is available for future work as needed.
criosWednesday, May 2, 2012 15:47:56On monady we were toldch B had bad tsys Guarionex and I double check all connection for losse screw. We re-alingned the 90 degrees connection on the filter.
crioscrios@naic.eduMonday, May 14, 2012 10:42:48Receiver tsys problem. The receiver was on load because the power cord of the antena/load power supply was broken.
Denis Urbaindurbain@naic.eduThursday, October 18, 2012 17:19:12The channel B had a much lower sensivity than Channel A. The Jennings Switch was temporalily replaced by an elbow and the sensitivity went back to a normal level. The Jennings switch was taken to the lab were we measured 6dB loss on the antenna position and 0.09dB on the load. The Jennings was adjusted with the help from Alfredo and the loss brought down to 0.1dB.
Dana Whitlowdwhitlow@naic.eduWednesday, December 26, 2012 15:11:34Today the RD/RA switch was replaced with a brand new unit. This was yet another attempt to understand/fix the now long-standing problem with Pol A's tsys being too high compared with that of Pol B. One test about two weeks ago had seemed to indicate that the then-installed switch was at fault, but now we're not so sure because the new switch does not seem to be helping. Next stop: bypassing the Jennings switch and other components in the path (yet again).
Denis Urbaindurbain@naic.eduFriday, January 25, 2013 16:39:49The pol B amplifier failed, both stages were damaged. Possible culprit the 430 transmiter from the carriage house because the Jennings switch had been removed and sent back for repair to the factory.The NRAO 450-12 was replaced by the 450-18 recently repaired by Richard Bradley in Charlottesville. Receiver pumping but not cooling yet.
Dana Whitlowdwhitlow@naic.eduMonday, February 4, 2013 15:09:24Disconnected Pol B receiver from feed and terminated the receiver's input, because of scheduled TX test later today and the absence of the Jennings switch (still out for repair) in Pol B.
Denis Urbaindurbain@naic.eduWednesday, April 3, 2013 13:22:53The Two Jennings switches came back from repair and one of them was installed back on the pol B side. The installation took place at the beginning of March after the switches were tested by Dana.
Denis Urbaindurbain@naic.eduFriday, August 9, 2013 16:41:59PolA has a higher Tsys than normal, the problem will be investigated next week. The waveguide assembly has been damaged during AO19 removal. The floor was rotated while the hoist was not stored away and the hook caught the wave-guide that has the adjustable back-short and bent it upwards. We will run some test to see if the performances have been affected and how serious is the damage.
Denis Urbaindurbain@naic.eduTuesday, September 3, 2013 15:12:07The mechanical damage of the turnstile was significant; the waveguide section that was grabbed by the hoist was bent outwards while the opposite side was bent inwards. The tilt of the all structure was such that the dual beam operation was impossible. The section of flexible waveguide could not be connected at both ends. The port isolation was measured after the accident and it was around 25dB which is not very good but could be consider acceptable. However since the dual beam operation was not possible, it was decided to replace the turnstile with the old one that was previously used. Only the termination arms are different, we removed the scattering post from the new turnstile and we had it installed in the new one. Some adapter plates were made in the machine shop to minimize the work done on the turnstile junction itself. The refurbished turnstile junction was installed and the isolation measured, about 30dB. This is better but not great. The monoplexer and switch assembly insertion loss was measured: 60dB which give overall a 90dB isolation between the transmitting arm and the receiving one. The Jennings switch was tested and replaced because of excessive insertion loss. The polA Tsys is now about 15K higher than the polB which is acceptable.
Dana Whitlowdwhitlow@naic.eduMonday, September 16, 2013 10:50:57Diagnosed new instabilities seen in 430GD receiver in the last several days (see http://www.naic.edu/~phil/cal430/cal430.html#13sep13). Problem seems clearly to be in the DCI BPF/notch filter (again!). We bypassed the filter and brought it down for another attempt at repair. With filter bypassed, the signal path seems solid as a rock, but users should anticipate problems with intermod distortion products from UHF TV stations for the duration. We'll get the filter fixed and put back in place ASAP.
Dana Whitlowdwhitlow@naic.eduTuesday, September 17, 2013 15:49:20The ailing bandpass filter was disassembled and all solder joints were redone. It then looked OK on the lab bench, and so was installed and placed back into service about 15:00 AST today (17 Sept 2013). Thanks to Guarionex and Anthony for the work they put into this repair effort.
Dana Whitlowdwhitlow@naic.eduFriday, January 3, 2014 16:23:34On Friday 27 Dec 2013, Chris Salter reported severe ripple in the Pol A passband of the 430 Gregorian receiver, and he submitted plots to prove his contention. These showed about 5 dB p-p ripple, a huge amount. The ripple period was about 3.1 MHz. Investigation did not initially give much of a clue, but it was noted that the amount of ripple was severely dependent on ZA. This led to a lot of effort spent in looking for a gravity-dependent effect on the stressed plumbing between the feed and the dewar's input, all to no avail. Also seriously considered and investigated was the possibility of a damaged optical fiber in the area where the cable drape suffered a problem a few weeks ago- again, no joy. Finally it was noticed that when the ZA was such as to make the ripple really severe (~10 dB p-p), one of the ripple peaks looked kind of narrow band and especially high, and upon closer inspection was seen to be band limited noise with a BW of ~50 kHz. So the theory was proposed that the problem was feedback and the system was barely shy of outright oscillation. The relevant cables on the receiver floor were carefully inspected, and finally it was discovered that the heliax SMA cable carrying RF from the post amp to the blue cabinet had a faulty shield connection at the connector. We've had this kind of trouble before with these cables on various receivers, but never with this bizarre set of symptoms. Upon repair of this cable, the problem went away and the passband looks very nice at all ZA settings.
Cold-head ratchetingfelix.fernandez@naic.edMonday, July 7, 2014 16:26:33Symptom: the receiver's cold-head was ratcheting. Procedure: It was decided to warm up the receiver, clean the cold-head's displacer and also clean the cold-head cylinder. The cold-head was replaced on July 07 2014 and the cool-down process was started.
Higher than normal Tsys on Pol-Afelix.fernandez@naic.eduMonday, August 4, 2014 13:46:27The power supply for the Jenning switches that control the ANT/LOAD state blew a fuse last Friday AUG-1-2014. The 327 receiver and Pol-A from the 430G receiver appeared to be warm during this past weekend’s Tsys. The fuse was replaced today and all systems are back to normal.
Cold-head contaminationfelix.fernandez@naic.eduMonday, August 11, 2014 12:02:20The 430G cold-head was ratcheting on AUG-07-2014. After removing the cold-head it was deemed that displacer was full of moisture. The displacer as "baked" and replaced. The cool-down process started today AUG-11-2014 at 11:00.
Col-head contemainationfelix.fernandez@naic.eduTuesday, October 21, 2014 15:10:12Once again, the cold-head was found to be ratcheting. A thermal cycle was started earlier today to service the cold-head. This contamination may be due to poor quality Helium. We will look into this a bit later.
Cold-head contaminationfelix.fernandez@naic.eduTuesday, October 28, 2014 15:40:52The cold-head was contaminated with humidity. The displacers were baked and the cold-head re-installed.
Post Amp Chassis Relay #4 Replacementfelix.fernandez@naic.eduTuesday, January 13, 2015 13:45:31Tsys readings were erratic on channel B during hcal, h90cal, lcal and l90cal. It was deemed that Realy #4 in the post-amp chassis was the culprit. Thus, Relay #was replaced and the problem resolved.
PolA Tsys readings unusually highfelix.fernandez@naic.eduFriday, August 14, 2015 09:58:07Since Tsys from AUG-12-2015 07:15 PolA has been showing Tsys values 100K (some times even more) above the expected values. The problem seems to have disappeared on it own since AUG-14-2015 07:15 Tsys. We checked for any problems with the system and the only thing that was done AFTER the problem disappeared was toggling the Radar/Antenna switch.
Sterling Cryo-Cooler Failure During Dual-Beamfelix.fernandez@naic.eduSaturday, September 16, 2017 17:18:49On SEP-16-2017 we noticed that the system temperature and physical temperature of the "warm" receiver used during dual-beam observations were high. We went up to the Dome and found that the sterling cryo-cooler was OFF. After a short inspection we found out that the 10A/32VP slow blow fuse on the cooler's power source was blown. After replacing the fuse the system started to work as expected. NOTES: A. The fuse housing became warm/hot to the touch once the new fuse was in place. We need to check what is causing this. B. The "Warm" receiver is actually a receiver cooled by a sterling cycle cryo-cooler which reaches temperature down to 90K. Why do we still call this "warm"?
Bad temperature cable connectorfelix.fernandez@naic.eduMonday, October 23, 2017 09:44:07The bias chassis' temperature cable connector has become loose. Temperature readings are always around 442K, which is typical when there is a missing connection. We brought down the bias chassis for further repairs.
Cold-head ratchtingfelix.fernandez@naic.eduFriday, November 3, 2017 09:47:58During the last week 430's cold head has been ratcheting. A purge procedure was completed Tuesday, but the system reverted to the previous ratcheting today (Friday). The system will be warmed up and the cold-head will be inspected next week.
Cold-head replacementfelix.fernandez@naic.eduTuesday, September 18, 2018 07:51:40Cold-head started ratcheting on 2018-09-18 (09:00) and a thermal cycle was started. Cold-head will be replaced.
Pol-B Jennings Relay Replacementfelix.fernandez@naic.eduTuesday, April 30, 2019 08:40:57The Tsys measurements have been showing erratic behavior on on Pol-B (every now and then it shows a Tsys about 20K higher than Pol-A). This is typically a faulty Jennings relay. Yesterday Pol-B Jennings Relay (SPDT) was replaced by one which was refurbished here at AO by A. Santoni.

Back to Top